# AlphaNet: Scaling Up Local Frame-based Atomistic Foundation Model

### v1

@@ Abstract
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com, haixiao@tsinghua.edu.cn and henkelman@utexas.edu.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [24, 3, 22]. While atomic forces necessary for these simulations can theoretically be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limit their application to small systems and short time scales. As a result, many phenomena occurring over longer time frames and larger spatial scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended periods [7, 23, 17]. However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
@@
The advent of machine learning has introduced a promising solution to this dilemma. By training models on data derived from ab initio calculations, machine learning forcefields (MLFF) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
@@
As interest in applying machine learning techniques to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 [35], is a large-scale pre-trained model with improved attention architecture over DeepPot [32]. Similarly, JMP [28] leverages diverse molecular systems of different types as a joint pre-training model. MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of machine learning-driven atomistic simulations [13, 2, 26, 29, 8].
However, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computation more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks which accuracy is more important.
@@
Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetry including rotation, translation and reflection. To reduce the unnecessary degree of freedoms, equivariant models are developed to build model that respects these symmetries  [25, 31]. Among all the developed models, most of the expressive models by date [2, 1, 21] are based on spherical harmonics [15]. However, the computation to calculate the tensor product of irreducible representations impose expensive computational overheads. While another branch of work which achieve rotation equivariance through building equivariant frames which could either be local or global [18, 9, 33, 12]. The main benefit of frame-based approaches is that is eliminates the necessity to use tensor product which greatly improves the computational efficiency.
@@
In this paper, we propose a local frame-based atomistic equivariant model (AlphaNet) for accurate yet highly efficient atomistic system simulations. Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate AlphaNet excels at accuracy, efficiency and scalability compared to state-of-the-art MLFF models on a variety of molecular systems, from defected graphene, formate decomposition, zeolites, to surface reactions.
@@ 2.2.2 Scaling
The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research. At present, our work has focused on scaling our model to the OC2M S2EF task, with plans to extend the training to the full OC20 dataset in future research. We compared the performance of our model against established models such as EquiformerV2 [21], SchNet [26], and DimeNet++ [14].
We were able to compare our results with SchNet and DimeNet++ models trained on the full OC20 dataset ( 134M data points), as well as with EquiformerV2 trained on both the full dataset and the smaller 2M subset. Currently we report our results on the validation set as it is public available. Results on the test set would be done as soon as possible. It’s important to note that each of these models was trained under different configurations, including training for both energy and force, training for energy only, and training for force only. The corresponding performance metrics are summarized in Table 2.
To control the training cost and make sure all of the elements are seen in the data, we first train our model on the full OC training set for 10,000 steps and then turns to 2M subset. We believe this would not provide additional benefits because we have tested that this procedure makes no difference in the validation ID set as there is no additional elements here. Our model trained on 2M subset for 2,000,000 steps demonstrates exceptional accuracy in energy prediction, achieving a mean absolute error (MAE) of 0.19 eV, which is significantly better than EquiformerV2’s performance on both the full dataset (0.229 eV) and the 2M subset (0.285 eV). SchNet and DimeNet++, which were trained on the full dataset, exhibit MAE values above 0.35 eV. This trend holds consistently across different test sets, demonstrating the robustness of our model in energy prediction.
In terms of force prediction, while our model trained on the 2M subset does not surpass larger models like EquiformerV2 and DimeNet++-large, it still achieves competitive results relative to models of a similar scale trained on the full dataset. Table 3 provides additional insights, revealing that our model has substantially lower training costs and fewer parameters compared to EquiformerV2 and DimeNet++. This efficiency, combined with the model’s high precision on this large dataset, indicates that our approach is computationally effective, requiring less data and resources while still delivering strong performance.
@@ 2.2.3 Speed
The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed. While increasing the number of parameters in an ML model can improve its precision, it often results in slower inference times. To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10. We also tested systems of varying sizes by expanding the unit cells. Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
We compared several graph-based models, including Nequip, MACE, SchNet, and PaiNN, against our AlphaNet. As illustrated in Figure 4, AlphaNet shows a distinct speed advantage over other MLFFs, with a relatively slow increase in computational time as system size grows. This demonstrates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
@@
Table 2: Energy and Force MAE for different models
Model ID OOD-adsorbate OOD-catalyst OOD-both
Energy MAE (eV) ↓
SchNet-full 0.43 0.49 0.53 0.72
SchNet-full-energy-only 0.38 0.45 0.55 0.70
SchNet-full-force-only 34.00 33.77 35.30 38.47
DimeNet++-full 0.48 0.47 0.53 0.65
DimeNet++-full-energy-only 0.35 0.40 0.51 0.65
DimeNet++-full-force-only 28.20 28.94 28.91 34.90
EquiformerV2-2M($\lambda_{E}=2$) 0.28 - - -
EquiformerV2-full($\lambda_{E}=2$) 0.22 - - -
AlphaNet-2M($\lambda_{E}=4$) 0.19 0.23 0.23 0.29
AlphaNet-2M($\lambda_{E}=20$) 0.18 0.23 0.25 0.30
Force MAE (eV/Å) ↓
SchNet-full 0.050 0.054 0.052 0.064
SchNet-full-energy-only 0.485 0.496 0.485 0.547
SchNet-full-force-only 0.045 0.048 0.047 0.058
DimeNet++-full 0.045 0.047 0.045 0.055
DimeNet++-full-energy-only 0.341 0.339 0.341 0.362
DimeNet++-full-force-only 0.034 0.035 0.035 0.041
EquiformerV2($\lambda_{E}=2$) 0.022 - - -
EquiformerV2-full($\lambda_{E}=2$) 0.016 - - -
AlphaNet-2M($\lambda_{E}=4$) 0.036 0.037 0.035 0.040
AlphaNet-2M($\lambda_{E}=20$) 0.054 0.062 0.061 0.070
Note: The downward arrow (↓) indicates that lower values are preferable. The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, ’2M’ indicates that the model was trained on OC2M sub set. $\lambda_{E}$ refers to the weight of the energy loss relative to the force, with a value set to 100. Bold indicates the best performance, while underline represents the second best.
Table 3: Training Performance and Training Time for Different Models
Models Energy MAE (eV) ↓ Force MAE (eV/Å) ↓ Training Time (GPU Days) ↓ Number of Parameters ↓
SchNet 0.549 0.0568 194 9.1M
DimeNet++ 0.515 0.0328 1600 10.1M
EquiformerV2 0.236 0.0157 1368 153M
AlphaNet 0.201 0.0356 360 6.1M
@@
Figure 4: Comparison of inference speed for various machine learning force fields. The x-axis represents the number of atoms in the systems, while the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches.

### v2

# AlphaNet: Scaling Up Local Frame-based Neural Interatomic Potential

@@ Abstract
00footnotetext: These authors contributed equally to this work.00footnotetext: Corresponding authors. Emails: wangjiaao0720@utexas.edu, yuanqidu@cs.cornell.edu, duanchenru@deepprinciple.com and haixiao@tsinghua.edu.cn.
@@ 1 Introduction
Molecular dynamics (MD) simulations have become essential for exploring and understanding complex phenomena in areas such as energy storage, catalysis, and biological systems [30, 4, 24]. While atomic forces necessary for these simulations can be derived from quantum mechanical approaches such as density functional theory (DFT), the computational cost associated with such first-principles methods severely limits their application to small-sized systems ($\sim{}10^{3}$ atoms) and short timescales ($\sim{}10^{1}$ ps). As a result, a wide spectrum of phenomena occurring over much larger scales remain inaccessible, even with the most powerful supercomputers.
@@
Classical force fields, which rely on predefined mathematical forms, offer a computationally efficient alternative, allowing simulations of larger systems over extended trajectories [8, 25, 19]. However, the simplicity of these models often compromises their accuracy, leading to a trade-off between computational speed and the fidelity of the simulated dynamics.
@@
The advent of machine learning techniques has introduced a promising solution to this dilemma. By training models on data derived from first-principles calculations, machine learning force fields (MLFFs) can potentially achieve the accuracy of first-principles methods while maintaining the computational efficiency of classical force fields. Unlike traditional models that rely on explicit functional forms to describe the bonded and non-bonded interactions, MLFFs are more flexible, learning to predict interactions based on the positions of atoms and their chemical identities.
@@
As interest in applying MLFFs to large-scale atomistic systems continues to grow, the field has seen a surge in the development of various innovative models. One example among the others, DPA-1 [43], is a large-scale pre-trained model with improved attention architecture over DeepPot [40]. Similarly, JMP [35] leverages diverse molecular systems of different types as a joint pre-training model. MACE [1], on the other hand, introduces higher-order message passing as a complete basis of many-body atomic interactions. Collectively, these models represent significant strides in the evolution of MLFF-driven atomistic simulations [14, 2, 33, 36, 9].
@@
Nevertheless, the balance between computational efficiency and accuracy is key to the practical application of these models. On one side, efficient yet less expressive models are more suitable for tasks that weigh computational cost more than accuracy. On the other side, less efficient yet expressive models are more suitable for tasks where accuracy is more important. Atomistic systems in 3D Euclidean space are invariant to Euclidean symmetries including rotation, translation and reflection. Thus, equivariant models are developed to respect these symmetries  [32, 38], and most of these expressive models by date [2, 1, 23] are based on the spherical harmonics [17], but calculating the tensor product of irreducible representations imposes expensive computational overheads. While another branch of work achieves rotation equivariance through building equivariant frames that can be either local or global [20, 10, 41, 13], and the main benefit of frame-based approaches is eliminating the necessity to involve the tensor product, thus greatly improving the computational efficiency.
@@
In this paper, we propose a local-frame-based atomistic equivariant model, named AlphaNet, for accurate yet highly efficient atomistic simulations. Building on the success of frame-based MLFFs in small atomistic systems, we introduce an additional rotary positional embedding to enhance the frame transition and temporal connection for multi-scale modeling. Extensive quantitative experiments demonstrate that AlphaNet excels at accuracy, efficiency and scalability compared with the state-of-the-art MLFF models on a variety of systems, ranging from defected graphenes, formate decomposition, zeolites, to surface reactions.
@@ 2.1 Dataset
The MPTraj Dataset and Matbench Discovery Test. MPTraj consists of 1,580,395 structures which are frames of the DFT relaxations performed on all 154,719 Materials Project materials [9]. Matbench Discovery [31] is a test task for predicting the stablity of 256,963 materials, which is also known as WBM test set [39]. In this case we train our model on the MPTraj dataset and test it on the WBM test set.
@@ 2.2.2 Scaling
The Open Catalyst Project (OCP) provides the OC20 dataset, a comprehensive collection of materials and structures designed to train machine learning models for catalysis research. In our current work, we focus on scaling our model to the OC2M S2EF task, with plans to extend training to the full OC20 dataset in future research. We compare the performance of our model against established benchmarks, including EquiformerV2 [23], EScAIP [29], eSCN [28], GemNet-OC [16], SchNet [33], and DimeNet++ [15].
Our comparisons include SchNet and DimeNet++ models trained on the full OC20 dataset ($\sim$134M data points), as well as EquiformerV2, EScAIP, eSCN, and GemNet-OC trained on the smaller 2M subset. Specifically, we utilize pretrained models of SchNet and DimeNet++ available at https://fair-chem.github.io/core/model_checkpoints.html, while results for the other models are sourced from [29]. All reported results are based on the validation set, as it is publicly accessible. The corresponding performance metrics are summarized in Table 2.
To manage training costs and ensure exposure to all elements in the dataset, we first train our model on the full OC training set for 10,000 steps before switching to the 2M subset. We note that this initial step does not provide additional benefits, as testing confirms no significant difference in performance on the validation ID set, which contains no additional elements. Our model, trained on the 2M subset for 2,000,000 steps, achieves exceptional accuracy in energy prediction, with a mean absolute error (MAE) of 0.24 eV. This performance is comparable to larger models like EquiformerV2 and EScAIP, while SchNet and DimeNet++, trained on the full dataset, exhibit higher MAE values above 0.35 eV. This trend remains consistent across various test sets, highlighting the robustness of our model in energy prediction.
In terms of force prediction, while our model trained on the 2M subset does not outperform larger models like EquiformerV2 and EScAIP, it remains competitive relative to models of similar scale trained on the full dataset. This may be attributed to the conservativeness of the models, which ensures forces are derived as $-\nabla_{x}E(x)$. Table 3 provides an efficiency evaluation of different models. Notably, EScAIP demonstrates remarkable efficiency in both speed and memory usage despite its large size, while AlphaNet shows a slight advantage in inference speed but consumes more memory than models like EScAIP. It is important to note that models such as AlphaNet, SchNet, and DimeNet++ are conservative, meaning their force calculations inherently require additional computational resources for autograd, resulting in higher time and memory costs.
Speak to WBM test, AlphaNet demonstrates a compelling balance between accuracy and physical consistency on the WBM benchmark 4. With an F1 score of 0.789 and DAF score of 4.315, it achieves near state-of-the-art classification performance while employing an EFSG-type conservative force field that rigorously enforces physical constraints. Notably, its compact architecture (16.2M parameters) outperforms the 31.2M-parameter eqV2 S DeNS model in regression tasks ($R^{2}=0.742$) with lower prediction error ($\text{MAE}=0.039$). This combination of efficiency and physics-aware design positions AlphaNet as a robust foundation for scalable materials modeling.
@@ 2.2.3 Speed
The primary advantage of machine learning force fields (MLFFs) over density functional theory (DFT) force fields lies in their significantly faster inference speed. While increasing the number of parameters in an ML model can enhance its precision, this often comes at the cost of slower inference times. To evaluate the speed of different MLFFs, we measured the average forward pass time across 2,000 zeolite structures with a batch size of 10. Additionally, we tested systems of varying sizes by expanding the unit cells. Although we aimed to maintain a consistent number of parameters across models, slight variations exist due to differences in architecture.
We compared several graph-based models, including NequIP, MACE, SchNet, and PaiNN, against our AlphaNet. As illustrated in Figure 4 (a), AlphaNet demonstrates a distinct speed advantage over other MLFFs, exhibiting a relatively slow increase in computational time as system size grows. This indicates that AlphaNet’s speed scales efficiently with larger systems, making it a more suitable option for handling increasing system sizes.
Furthermore, we evaluated the inference speed of pretrained models on the Matbench Discovery leaderboard. To ensure fairness, we selected models that are compliant, open-source, and accessible without additional requests. We measured the average speed of energy prediction using the same API. The models included AlphaNet, ORB-v2 [26], SevenNet [27], GRACE [3], MACE, and CHGNet. Figure 4 (b) presents the results. Among these models, GRACE, which is based on TensorFlow, exhibits remarkable speed compared to the others, which are PyTorch-based. Notably, AlphaNet-MPtrj’s inference speed is comparable to models with 10 times fewer parameters (e.g., SevenNet-0 and SevenNet-l3i5). This suggests that AlphaNet achieves competitive efficiency despite its larger parameter count, and it would be intriguing to explore its performance with fewer parameters in future work.
@@
Table 2: Results on the OC20 validation set
Model Parameters ↓ Energy MAE (eV) ↓ Force MAE (eV/Å) ↓
SchNet-full 9.1M 0.54 0.55
DimeNet++-full 10.1M 0.53 0.048
GemNet-OC 38M 0.29 0.026
eSCN 51M 0.28 0.021
EScAIP-Medium 146M 0.25 0.019
EquiformerV2($\lambda_{E}=2$) 146M 0.28 0.022
AlphaNet($\lambda_{E}=4$) 6.1M 0.25 0.040
AlphaNet($\lambda_{E}=20$) 6.1M 0.24 0.062
Note: The downward arrow (↓) indicates that lower values are preferable. The term ’full’ in a model’s name signifies that it was trained on the complete OC20 training set, the others are trained on the 2M subset. $\lambda_{E}$ refers to the weight of the energy loss relative to the force, with a value set to 100. Bold indicates the best performance.
Table 3: Efficiency on OC dataset, measured on A100 40GB GPU
Models Parameters Training Speed (Sample/Sec) ↑ Training Memory (GB/Sample) ↓ Inference Speed (Sample/Sec) ↑ Inference Memory (GB/Sample) ↓ Conserv.
SchNet 9.1M 39 2.5 112.1 1.7
DimeNet++ 10.1M 29 3.6 92 2.5
GemNet-OC 38M 16.22 1.63 39.2 1.61 $\times$
eSCN 51M 8.5 1.75 25.2 1.74 $\times$
EScAIP 83M 37.8 1.23 111.8 1.09 $\times$
EquiformerV2 31M 15.7 1.63 36.72 1.61 $\times$
AlphaNet 6.1M 41.2 2.2 120.1 1.5 $\checkmark$
Table 4: Results on WBM full test set
Model F1 ↑ DAF ↑ Prec ↑ Acc ↑ MAE ↓ R2 ↑ Size Type
eqV2 S DeNS 0.798 4.362 0.748 0.927 0.035 0.785 31.2M EFSD
AlphaNet (ours) 0.789 4.315 0.740 0.923 0.039 0.742 16.2M EFSG
DPA3-v1-MPtrj 0.757 4.13 0.709 0.911 0.04 0.793 3.37M EFSG
ORB MPtrj 0.755 4.188 0.719 0.911 0.043 0.752 25.2M EFSD
SevenNet-l3i5 0.751 4.112 0.706 0.909 0.042 0.773 1.17M EFSG
SevenNet-0 0.719 3.804 0.653 0.893 0.046 0.75 842k EFSG
GRACE-2L-MPtrj 0.687 3.714 0.637 0.884 0.05 0.74 15.3M EFSG
MACE-MP-0 0.668 3.4 0.583 0.867 0.055 0.698 4.69M EFSG
CHGNet 0.612 3.038 0.521 0.839 0.061 0.69 413k EFSGM
 [^1] Data source:Open Catalyst https://fair-chem.github.io/core/model_checkpoints.html; Matbench Discovery https://matbench-discovery.materialsproject.org/
@@
Figure 4: Comparison of inference speed for various machine learning force fields. (a) Inference speed as a function of system size. The x-axis represents the number of atoms in the system, and the y-axis shows the average time required to predict energy and forces for systems in batches of 10, across a total of 200 batches. (b) Inference speed and F1 performance of models accessible and compliant with the Matbench Discovery leaderboard. The speed is calculated as the number of energies predicted per second.
@@ 3 Discussion
Improving the smoothness of the output is also crucial, as we believe it would enhance the models’ generalization ability and yield a more physically meaningful potential energy surface.